Musich Corporation has an activity-based costing system with three activity cost pools--Machining, Setting Up, and Other. The company's overhead costs, which consist of equipment depreciation and indirect labor, have been allocated to the cost pools already and are provided in the table below.

                                                               Activity Cost Pools

Machining

Setting Up

Other

Total

Equipment depreciation

$

9200

$

48,200

$

23,600

$

81,000

Indirect labor

4200

2800

4000

11,000

Total

$

13,400

$

51,000

$

27,600

$

92,000

Costs in the Machining cost pool are assigned to products based on machine-hours (MHs) and costs in the Setting Up cost pool are assigned to products based on the number of batches. Costs in the Other cost pool are not assigned to products. Data concerning the two products and the company's costs appear below:

MHs

Batches

Product Z3

5900

700

Product T1

6100

1300

Total

12,000

2000

                                                                                         Product Z3      Product T1

Sales (total)

$

228,800

$

254,500

Direct materials (total)

$

83,500

$

97,400

Direct labor (total)

$

109,800

$

104,700

Required:

Calculate the following:

Machining Activity Rate

$

Setting up Activity Rate

$

Amount of OH applied to product Z3 (round to the nearest dollar)

$

Amount of OH applied to product T1 (round to the nearest dollar)

$

Product Margin – Z3 (round to the nearest dollar)

$

Product Margin – T1 (round to the nearest dollar)

$
Answer #1

1.

Machining activity rate = Total overhead / Total machine hours

Machining activity rate = $13,400 / 12,000 = $1.12

Setting up activity rate = Total overhead / Total batches

Setting up activity rate = $51,000 / 2,000 = $25.5 per batch

2.

Overhead applied to Product Z3 = 5,900*$1.12 + 700*$25.5

Overhead applied to Product Z3 = $6,608 + 17,850 = $24,458

Overhead applied to Product T1 = 6,100*$1.12 + 1,300*$25.5

Overhead applied to Product T1 = $6,832 + 33,150 = $39,982

3.

Product Z3 Product T1
Sales $228,800 $254,500
Less: Direct material 83,500 97,400
Direct labor 109,800 104,700
Overhead applied 24,458 39,982
Product margin $11,042 $12,418